9. Determine the energy of an electromagnetic radiation of 550 nm wavelength.
Answer #1

Given:

lambda = 550 nm = 5.50*10^-7 m

use:

E = h*c/lambda

=(6.626*10^-34 J.s)*(3.0*10^8 m/s)/(5.5*10^-7 m)

= 3.614*10^-19 J

Answer: 3.61*10^-19 J